# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ Appendix F Algorithms
mix box_count_threshold thresholded lac additive 15.4120 0.0219 7.1302 0.0932 10.0158 0.0960 0.1645
mix box_count_threshold thresholded lac multiplicative 15.4120 0.0219 4.1099 0.1008 10.0158 0.0960 0.1797
mix box_count_threshold thresholded aps additive 15.4120 0.0219 7.1302 0.0932 10.2455 0.0945 0.1645
mix box_count_threshold thresholded aps multiplicative 15.4120 0.0219 3.8031 0.1008 8.4829 0.0960 0.1797
mix box_count_threshold pixelwise lac additive 15.4120 0.0219 2.9879 0.0992 0.5591 0.1045 0.1797
mix box_count_threshold pixelwise lac multiplicative 15.4120 0.0219 1.7427 0.1024 0.5591 0.1045 0.1830
mix box_count_threshold pixelwise aps additive 15.4120 0.0219 2.9879 0.0992 0.9931 0.0611 0.1449
mix box_count_threshold pixelwise aps multiplicative 15.4120 0.0219 1.7427 0.1024 0.9817 0.0611 0.1473
mix box_count_threshold boxwise lac additive 15.4120 0.0219 3.4491 0.0992 0.5591 0.1045 0.1797
mix box_count_threshold boxwise lac multiplicative 15.4120 0.0219 2.2983 0.1024 0.5591 0.1045 0.1830
mix box_count_threshold boxwise aps additive 15.4120 0.0219 3.4491 0.0992 0.9931 0.0611 0.1449
mix box_count_threshold boxwise aps multiplicative 15.4120 0.0219 2.2983 0.1024 0.9817 0.0611 0.1473
mix box_count_recall thresholded lac additive 10.1384 0.0310 7.1302 0.0932 10.0158 0.0960 0.1645
mix box_count_recall thresholded lac multiplicative 10.1384 0.0310 4.1099 0.1008 10.0158 0.0960 0.1797
mix box_count_recall thresholded aps additive 10.1384 0.0310 7.1302 0.0932 10.2455 0.0945 0.1645
mix box_count_recall thresholded aps multiplicative 10.1384 0.0310 3.8031 0.1008 8.4829 0.0960 0.1797
mix box_count_recall pixelwise lac additive 10.1384 0.0310 2.9879 0.0992 0.5591 0.1045 0.1797
mix box_count_recall pixelwise lac multiplicative 10.1384 0.0310 1.7427 0.1024 0.5591 0.1045 0.1830
mix box_count_recall pixelwise aps additive 10.1384 0.0310 2.9879 0.0992 0.9931 0.0611 0.1449
mix box_count_recall pixelwise aps multiplicative 10.1384 0.0310 1.7427 0.1024 0.9817 0.0611 0.1473
mix box_count_recall boxwise lac additive 10.1384 0.0310 3.4491 0.0992 0.5591 0.1045 0.1797
mix box_count_recall boxwise lac multiplicative 10.1384 0.0310 2.2983 0.1024 0.5591 0.1045 0.1830
mix box_count_recall boxwise aps additive 10.1384 0.0310 3.4491 0.0992 0.9931 0.0611 0.1449
mix box_count_recall boxwise aps multiplicative 10.1384 0.0310 2.2983 0.1024 0.9817 0.0611 0.1473
giou box_count_threshold thresholded lac additive 15.4120 0.0219 19.5331 0.0960 26.8823 0.0968 0.1684
giou box_count_threshold thresholded lac multiplicative 15.4120 0.0219 14.1941 0.1000 26.8823 0.0968 0.1797
giou box_count_threshold thresholded aps additive 15.4120 0.0219 19.5331 0.0960 26.6024 0.0953 0.1676
giou box_count_threshold thresholded aps multiplicative 15.4120 0.0219 13.7640 0.1000 22.8826 0.0968 0.1797
giou box_count_threshold pixelwise lac additive 15.4120 0.0219 16.3220 0.0984 26.8823 0.0968 0.1712
giou box_count_threshold pixelwise lac multiplicative 15.4120 0.0219 13.3474 0.1008 26.8823 0.0968 0.1797
giou box_count_threshold pixelwise aps additive 15.4120 0.0219 16.3220 0.0984 26.6024 0.0953 0.1676
giou box_count_threshold pixelwise aps multiplicative 15.4120 0.0219 13.1761 0.1008 22.8826 0.0968 0.1797
giou box_count_threshold boxwise lac additive 15.4120 0.0219 17.0479 0.0984 26.8823 0.0968 0.1712
giou box_count_threshold boxwise lac multiplicative 15.4120 0.0219 14.1941 0.1000 26.8823 0.0968 0.1797
giou box_count_threshold boxwise aps additive 15.4120 0.0219 17.0479 0.0984 26.6024 0.0953 0.1676
giou box_count_threshold boxwise aps multiplicative 15.4120 0.0219 13.7640 0.1000 22.8826 0.0968 0.1797
giou box_count_recall thresholded lac additive 10.1384 0.0310 19.5331 0.0960 26.8823 0.0968 0.1684
giou box_count_recall thresholded lac multiplicative 10.1384 0.0310 14.1941 0.1000 26.8823 0.0968 0.1797
giou box_count_recall thresholded aps additive 10.1384 0.0310 19.5331 0.0960 26.6024 0.0953 0.1676
giou box_count_recall thresholded aps multiplicative 10.1384 0.0310 13.7640 0.1000 22.8826 0.0968 0.1797
giou box_count_recall pixelwise lac additive 10.1384 0.0310 16.3220 0.0984 26.8823 0.0968 0.1712
giou box_count_recall pixelwise lac multiplicative 10.1384 0.0310 13.3474 0.1008 26.8823 0.0968 0.1797
giou box_count_recall pixelwise aps additive 10.1384 0.0310 16.3220 0.0984 26.6024 0.0953 0.1676
giou box_count_recall pixelwise aps multiplicative 10.1384 0.0310 13.1761 0.1008 22.8826 0.0968 0.1797
giou box_count_recall boxwise lac additive 10.1384 0.0310 17.0479 0.0984 26.8823 0.0968 0.1712
giou box_count_recall boxwise lac multiplicative 10.1384 0.0310 14.1941 0.1000 26.8823 0.0968 0.1797
giou box_count_recall boxwise aps additive 10.1384 0.0310 17.0479 0.0984 26.6024 0.0953 0.1676
giou box_count_recall boxwise aps multiplicative 10.1384 0.0310 13.7640 0.1000 22.8826 0.0968 0.1797
hausdorff box_count_threshold thresholded lac additive 15.4120 0.0219 4.7787 0.0901 25.4892 0.0984 0.1657
hausdorff box_count_threshold thresholded lac multiplicative 15.4120 0.0219 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_threshold thresholded aps additive 15.4120 0.0219 4.7787 0.0901 24.3324 0.0960 0.1657
hausdorff box_count_threshold thresholded aps multiplicative 15.4120 0.0219 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_threshold pixelwise lac additive 15.4120 0.0219 2.9569 0.1008 25.4892 0.0984 0.1797
hausdorff box_count_threshold pixelwise lac multiplicative 15.4120 0.0219 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_threshold pixelwise aps additive 15.4120 0.0219 2.9569 0.1008 24.3324 0.0960 0.1765
hausdorff box_count_threshold pixelwise aps multiplicative 15.4120 0.0219 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_threshold boxwise lac additive 15.4120 0.0219 3.9352 0.0960 25.4892 0.0984 0.1692
hausdorff box_count_threshold boxwise lac multiplicative 15.4120 0.0219 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_threshold boxwise aps additive 15.4120 0.0219 3.9352 0.0960 24.3324 0.0960 0.1692
hausdorff box_count_threshold boxwise aps multiplicative 15.4120 0.0219 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_recall thresholded lac additive 10.1384 0.0310 4.7787 0.0901 25.4892 0.0984 0.1657
hausdorff box_count_recall thresholded lac multiplicative 10.1384 0.0310 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_recall thresholded aps additive 10.1384 0.0310 4.7787 0.0901 24.3324 0.0960 0.1657
hausdorff box_count_recall thresholded aps multiplicative 10.1384 0.0310 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_recall pixelwise lac additive 10.1384 0.0310 2.9569 0.1008 25.4892 0.0984 0.1797
hausdorff box_count_recall pixelwise lac multiplicative 10.1384 0.0310 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_recall pixelwise aps additive 10.1384 0.0310 2.9569 0.1008 24.3324 0.0960 0.1765
hausdorff box_count_recall pixelwise aps multiplicative 10.1384 0.0310 2.5616 0.1000 24.3324 0.0960 0.1797
hausdorff box_count_recall boxwise lac additive 10.1384 0.0310 3.9352 0.0960 25.4892 0.0984 0.1692
hausdorff box_count_recall boxwise lac multiplicative 10.1384 0.0310 2.5616 0.1000 25.4892 0.0984 0.1797
hausdorff box_count_recall boxwise aps additive 10.1384 0.0310 3.9352 0.0960 24.3324 0.0960 0.1692
hausdorff box_count_recall boxwise aps multiplicative 10.1384 0.0310 2.5616 0.1000 24.3324 0.0960 0.1797
lac box_count_threshold thresholded lac additive 15.4120 0.0219 12.7995 0.0984 0.5591 0.1045 0.1757
lac box_count_threshold thresholded lac multiplicative 15.4120 0.0219 8.1480 0.1008 0.5591 0.1045 0.1797
lac box_count_threshold thresholded aps additive 15.4120 0.0219 12.7995 0.0984 0.9931 0.0611 0.1449
lac box_count_threshold thresholded aps multiplicative 15.4120 0.0219 8.1480 0.1008 0.9817 0.0611 0.1473
lac box_count_threshold pixelwise lac additive 15.4120 0.0219 9.6219 0.1000 0.5591 0.1045 0.1797
lac box_count_threshold pixelwise lac multiplicative 15.4120 0.0219 6.8758 0.1016 0.5591 0.1045 0.1830
lac box_count_threshold pixelwise aps additive 15.4120 0.0219 9.6219 0.1000 0.9931 0.0611 0.1473
lac box_count_threshold pixelwise aps multiplicative 15.4120 0.0219 6.8758 0.1016 0.9817 0.0611 0.1473
lac box_count_threshold boxwise lac additive 15.4120 0.0219 10.6310 0.1000 0.5591 0.1045 0.1797
lac box_count_threshold boxwise lac multiplicative 15.4120 0.0219 8.1480 0.1008 0.5591 0.1045 0.1797
lac box_count_threshold boxwise aps additive 15.4120 0.0219 10.6310 0.1000 0.9931 0.0611 0.1473
lac box_count_threshold boxwise aps multiplicative 15.4120 0.0219 8.1480 0.1008 0.9817 0.0611 0.1473
lac box_count_recall thresholded lac additive 10.1384 0.0310 12.7995 0.0984 0.5591 0.1045 0.1757
lac box_count_recall thresholded lac multiplicative 10.1384 0.0310 8.1480 0.1008 0.5591 0.1045 0.1797
lac box_count_recall thresholded aps additive 10.1384 0.0310 12.7995 0.0984 0.9931 0.0611 0.1449
lac box_count_recall thresholded aps multiplicative 10.1384 0.0310 8.1480 0.1008 0.9817 0.0611 0.1473
lac box_count_recall pixelwise lac additive 10.1384 0.0310 9.6219 0.1000 0.5591 0.1045 0.1797
lac box_count_recall pixelwise lac multiplicative 10.1384 0.0310 6.8758 0.1016 0.5591 0.1045 0.1830
lac box_count_recall pixelwise aps additive 10.1384 0.0310 9.6219 0.1000 0.9931 0.0611 0.1473
lac box_count_recall pixelwise aps multiplicative 10.1384 0.0310 6.8758 0.1016 0.9817 0.0611 0.1473
lac box_count_recall boxwise lac additive 10.1384 0.0310 10.6310 0.1000 0.5591 0.1045 0.1797
lac box_count_recall boxwise lac multiplicative 10.1384 0.0310 8.1480 0.1008 0.5591 0.1045 0.1797
lac box_count_recall boxwise aps additive 10.1384 0.0310 10.6310 0.1000 0.9931 0.0611 0.1473
lac box_count_recall boxwise aps multiplicative 10.1384 0.0310 8.1480 0.1008 0.9817 0.0611 0.1473

### v2

@@ Appendix F Effect of Object Detection and Conformal Settings
mix box_count_threshold thresholded lac additive 15.7624 0.0212 5.6266 0.0928 57.0939 0.1017 0.1811
mix box_count_threshold thresholded lac multiplicative 15.7624 0.0212 4.8604 0.0944 57.0939 0.1017 0.1916
mix box_count_threshold thresholded aps additive 15.7624 0.0212 5.6266 0.0928 22.5433 0.0856 0.169
mix box_count_threshold thresholded aps multiplicative 15.7624 0.0212 4.8604 0.0944 22.5433 0.0856 0.1756
mix box_count_threshold pixelwise lac additive 15.7624 0.0212 3.2599 0.1012 57.0939 0.1017 0.1753
mix box_count_threshold pixelwise lac multiplicative 15.7624 0.0212 2.5269 0.1009 57.0939 0.1017 0.1824
mix box_count_threshold pixelwise aps additive 15.7624 0.0212 3.2599 0.1012 22.5433 0.0856 0.1626
mix box_count_threshold pixelwise aps multiplicative 15.7624 0.0212 2.5269 0.1009 22.5433 0.0856 0.1691
mix box_count_threshold boxwise lac additive 15.7624 0.0212 5.4937 0.0945 57.0939 0.1017 0.1717
mix box_count_threshold boxwise lac multiplicative 15.7624 0.0212 3.8395 0.0988 57.0939 0.1017 0.1825
mix box_count_threshold boxwise aps additive 15.7624 0.0212 5.4937 0.0945 22.5433 0.0856 0.1593
mix box_count_threshold boxwise aps multiplicative 15.7624 0.0212 3.8395 0.0988 22.5433 0.0856 0.1686
mix box_count_recall thresholded lac additive 10.1584 0.0313 4.7828 0.0896 54.0995 0.0994 0.1754
mix box_count_recall thresholded lac multiplicative 10.1584 0.0313 5.1396 0.0936 54.0995 0.0994 0.1883
mix box_count_recall thresholded aps additive 10.1584 0.0313 4.7828 0.0896 24.5079 0.0818 0.16
mix box_count_recall thresholded aps multiplicative 10.1584 0.0313 5.1396 0.0936 24.5079 0.0818 0.1695
mix box_count_recall pixelwise lac additive 10.1584 0.0313 2.9593 0.1007 54.0995 0.0994 0.172
mix box_count_recall pixelwise lac multiplicative 10.1584 0.0313 2.6906 0.0984 54.0995 0.0994 0.1787
mix box_count_recall pixelwise aps additive 10.1584 0.0313 2.9593 0.1007 24.5079 0.0818 0.1562
mix box_count_recall pixelwise aps multiplicative 10.1584 0.0313 2.6906 0.0984 24.5079 0.0818 0.1623
mix box_count_recall boxwise lac additive 10.1584 0.0313 4.7128 0.0904 54.0995 0.0994 0.1645
mix box_count_recall boxwise lac multiplicative 10.1584 0.0313 4.0727 0.0942 54.0995 0.0994 0.1767
mix box_count_recall boxwise aps additive 10.1584 0.0313 4.7128 0.0904 24.5079 0.0818 0.1497
mix box_count_recall boxwise aps multiplicative 10.1584 0.0313 4.0727 0.0942 24.5079 0.0818 0.1596
hausdorff box_count_threshold thresholded lac additive 15.7624 0.0212 5.6266 0.0908 58.4583 0.1018 0.1803
hausdorff box_count_threshold thresholded lac multiplicative 15.7624 0.0212 5.3222 0.0892 58.4583 0.1018 0.1866
hausdorff box_count_threshold thresholded aps additive 15.7624 0.0212 5.6266 0.0908 23.2279 0.0871 0.1684
hausdorff box_count_threshold thresholded aps multiplicative 15.7624 0.0212 5.3222 0.0892 23.2279 0.0871 0.1724
hausdorff box_count_threshold pixelwise lac additive 15.7624 0.0212 3.2677 0.1007 58.4583 0.1018 0.1752
hausdorff box_count_threshold pixelwise lac multiplicative 15.7624 0.0212 2.6485 0.0966 58.4583 0.1018 0.1797
hausdorff box_count_threshold pixelwise aps additive 15.7624 0.0212 3.2677 0.1007 23.2279 0.0871 0.1632
hausdorff box_count_threshold pixelwise aps multiplicative 15.7624 0.0212 2.6485 0.0966 23.2279 0.0871 0.167
hausdorff box_count_threshold boxwise lac additive 15.7624 0.0212 5.4858 0.0953 58.4583 0.1018 0.1724
hausdorff box_count_threshold boxwise lac multiplicative 15.7624 0.0212 4.1312 0.0918 58.4583 0.1018 0.1768
hausdorff box_count_threshold boxwise aps additive 15.7624 0.0212 5.4858 0.0953 23.2279 0.0871 0.1612
hausdorff box_count_threshold boxwise aps multiplicative 15.7624 0.0212 4.1312 0.0918 23.2279 0.0871 0.164
hausdorff box_count_recall thresholded lac additive 10.1584 0.0313 4.7828 0.088 55.4691 0.0992 0.1739
hausdorff box_count_recall thresholded lac multiplicative 10.1584 0.0313 5.6488 0.0824 55.4691 0.0992 0.1777
hausdorff box_count_recall thresholded aps additive 10.1584 0.0313 4.7828 0.088 24.8851 0.0835 0.16
hausdorff box_count_recall thresholded aps multiplicative 10.1584 0.0313 5.6488 0.0824 24.885 0.0835 0.1608
hausdorff box_count_recall pixelwise lac additive 10.1584 0.0313 2.9721 0.0996 55.4692 0.0992 0.171
hausdorff box_count_recall pixelwise lac multiplicative 10.1584 0.0313 2.8361 0.0924 55.4692 0.0992 0.1739
hausdorff box_count_recall pixelwise aps additive 10.1584 0.0313 2.9721 0.0996 24.8851 0.0835 0.1566
hausdorff box_count_recall pixelwise aps multiplicative 10.1584 0.0313 2.8361 0.0924 24.885 0.0835 0.1589
hausdorff box_count_recall boxwise lac additive 10.1584 0.0313 4.7128 0.0898 55.4691 0.0992 0.1642
hausdorff box_count_recall boxwise lac multiplicative 10.1584 0.0313 4.3879 0.0845 55.4691 0.0992 0.1685
hausdorff box_count_recall boxwise aps additive 10.1584 0.0313 4.7128 0.0898 24.8851 0.0835 0.1505
hausdorff box_count_recall boxwise aps multiplicative 10.1584 0.0313 4.3879 0.0845 24.885 0.0835 0.1532
lac box_count_threshold thresholded lac additive 15.7624 0.0212 19.6105 0.09 0.5361 0.106 0.1837
lac box_count_threshold thresholded lac multiplicative 15.7624 0.0212 13.3677 0.1008 0.5361 0.106 0.1881
lac box_count_threshold thresholded aps additive 15.7624 0.0212 19.6105 0.09 0.9956 0.0461 0.1288
lac box_count_threshold thresholded aps multiplicative 15.7624 0.0212 13.3677 0.1008 0.9956 0.0461 0.1378
lac box_count_threshold pixelwise lac additive 15.7624 0.0212 14.1831 0.0913 0.5361 0.106 0.1686
lac box_count_threshold pixelwise lac multiplicative 15.7624 0.0212 8.3605 0.1051 0.5361 0.106 0.174
lac box_count_threshold pixelwise aps additive 15.7624 0.0212 14.1831 0.0913 0.9956 0.0461 0.1218
lac box_count_threshold pixelwise aps multiplicative 15.7624 0.0212 8.3605 0.1051 0.9956 0.0461 0.133
lac box_count_threshold boxwise lac additive 15.7624 0.0212 17.3806 0.0919 0.5361 0.106 0.1667
lac box_count_threshold boxwise lac multiplicative 15.7624 0.0212 9.7946 0.1068 0.5361 0.106 0.1728
lac box_count_threshold boxwise aps additive 15.7624 0.0212 17.3806 0.0919 0.9956 0.0461 0.1203
lac box_count_threshold boxwise aps multiplicative 15.7624 0.0212 9.7946 0.1068 0.9956 0.0461 0.1324
lac box_count_recall thresholded lac additive 10.1584 0.0313 16.6423 0.0844 0.7161 0.106 0.1788
lac box_count_recall thresholded lac multiplicative 10.1584 0.0313 14.4994 0.0904 0.7161 0.106 0.1794
lac box_count_recall thresholded aps additive 10.1584 0.0313 16.6423 0.0844 0.9932 0.0611 0.1369
lac box_count_recall thresholded aps multiplicative 10.1584 0.0313 14.4994 0.0904 0.9932 0.0611 0.1414
lac box_count_recall pixelwise lac additive 10.1584 0.0313 12.236 0.0859 0.7161 0.106 0.1631
lac box_count_recall pixelwise lac multiplicative 10.1584 0.0313 8.9708 0.0974 0.7161 0.106 0.1675
lac box_count_recall pixelwise aps additive 10.1584 0.0313 12.236 0.0859 0.9932 0.0611 0.1263
lac box_count_recall pixelwise aps multiplicative 10.1584 0.0313 8.9708 0.0974 0.9932 0.0611 0.1357
lac box_count_recall boxwise lac additive 10.1584 0.0313 14.9561 0.0845 0.7161 0.106 0.1608
lac box_count_recall boxwise lac multiplicative 10.1584 0.0313 10.5469 0.0966 0.7161 0.106 0.1659
lac box_count_recall boxwise aps additive 10.1584 0.0313 14.9561 0.0845 0.9932 0.0611 0.1238
lac box_count_recall boxwise aps multiplicative 10.1584 0.0313 10.5469 0.0966 0.9932 0.0611 0.1338
giou box_count_threshold thresholded lac additive 15.7624 0.0212 39.4871 0.0348 57.1519 0.0891 0.123
giou box_count_threshold thresholded lac multiplicative 15.7624 0.0212 179.9897 0.0196 57.1519 0.0891 0.1084
giou box_count_threshold thresholded aps additive 15.7624 0.0212 39.4871 0.0348 57.7448 0.0459 0.0792
giou box_count_threshold thresholded aps multiplicative 15.7624 0.0212 179.9897 0.0196 57.7448 0.0459 0.0651
giou box_count_threshold pixelwise lac additive 15.7624 0.0212 34.0626 0.0495 57.1519 0.0891 0.1346
giou box_count_threshold pixelwise lac multiplicative 15.7624 0.0212 140.1268 0.0222 57.1519 0.0891 0.11
giou box_count_threshold pixelwise aps additive 15.7624 0.0212 34.0626 0.0495 57.7448 0.0459 0.0896
giou box_count_threshold pixelwise aps multiplicative 15.7624 0.0212 140.1268 0.0222 57.7448 0.0459 0.0667
giou box_count_threshold boxwise lac additive 15.7624 0.0212 38.7389 0.0397 57.1519 0.0891 0.1257
giou box_count_threshold boxwise lac multiplicative 15.7624 0.0212 161.3222 0.0192 57.1519 0.0891 0.1072
giou box_count_threshold boxwise aps additive 15.7624 0.0212 38.7389 0.0397 57.7448 0.0459 0.0803
giou box_count_threshold boxwise aps multiplicative 15.7624 0.0212 161.3222 0.0192 57.7448 0.0459 0.0639
giou box_count_recall thresholded lac additive 10.1584 0.0313 32.3011 0.022 58.219 0.0858 0.1066
giou box_count_recall thresholded lac multiplicative 10.1584 0.0313 179.9438 0.0108 58.219 0.0858 0.096
giou box_count_recall thresholded aps additive 10.1584 0.0313 32.3011 0.022 58.5988 0.0432 0.064
giou box_count_recall thresholded aps multiplicative 10.1584 0.0313 179.9438 0.0108 58.5988 0.0432 0.0534
giou box_count_recall pixelwise lac additive 10.1584 0.0313 27.9418 0.0375 58.219 0.0858 0.121
giou box_count_recall pixelwise lac multiplicative 10.1584 0.0313 140.4195 0.0123 58.219 0.0858 0.0974
giou box_count_recall pixelwise aps additive 10.1584 0.0313 27.9418 0.0375 58.5988 0.0432 0.0759
giou box_count_recall pixelwise aps multiplicative 10.1584 0.0313 140.4195 0.0123 58.5988 0.0432 0.0546
giou box_count_recall boxwise lac additive 10.1584 0.0313 31.7182 0.0274 58.2189 0.0858 0.1108
giou box_count_recall boxwise lac multiplicative 10.1584 0.0313 162.1215 0.0113 58.2189 0.0858 0.0964
giou box_count_recall boxwise aps additive 10.1584 0.0313 31.7182 0.0274 58.5988 0.0432 0.0665
giou box_count_recall boxwise aps multiplicative 10.1584 0.0313 162.1215 0.0113 58.5988 0.0432 0.0536